# SpotLight: Shadow-Guided Object Relighting via Diffusion

### v1

@@ Abstract
Recent work has shown that diffusion models can be used as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. Unlike typical physics-based renderers, however, neural rendering engines are limited by the lack of manual control over the lighting setup, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise lighting control can be achieved for object relighting simply by specifying the desired shadows of the object. Rather surprisingly, we show that injecting only the shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training. Specifically, we demonstrate its use with two neural renderers from the recent literature. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting.
| Unconditioned | Shadow-conditioned (top: shadow mapping, bottom: hand-drawn scribbles) composites from SpotLight | |||||
@@ 1 Introduction
Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process. When one is given a single photograph, however, no such control is possible. Relighting then becomes a hard task because of the complex surface interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture. The recent rise of powerful image generation models (e.g., [52]) has enabled exciting possibilities for image relighting [32, 73].
@@
This paper focuses on providing local lighting control, where one wishes to modify the appearance of a single object. This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image. In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
In this paper, we present SpotLight, a method for realistic, controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than training a model specifically for the task of object relighting, our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow to a general-purpose neural renderer. Indeed, we find that simply specifying the desired lighting through the target shadow makes the renderer relight the object accordingly! SpotLight: Shadow-Guided Object Relighting via Diffusion shows two examples of such shadow-based lighting control: 1) the shadow of the couch is rendered using basic shadow mapping [19] according to the desired direction (top); and 2) one can even draw scribbles by hand to depict the desired shadow of the ottoman (bottom). To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained neural renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
@@
The core technical contribution lies in our shadow injection approach which is robust to artifacts in the target shadow map. We build upon existing object compositing networks [75, 72] that operate on intrinsic images and exploit diffusion models’ learned priors to render realistic outputs. We demonstrate SpotLight handles complex scenarios and maintains physical plausibility without requiring extensive manual intervention nor any additional training.
@@
In summary, our contributions are as follows. First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring fine-tuning or additional training. We demonstrate superior quality, both perceptually through a user study and quantitatively, compared to competitive specialized models. Finally, we release a novel evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
@@ 2 Related work
Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [58, 28, 59, 52]. These models offer control mechanisms for various modalities, such as text [48, 52], image content [53, 21], intrinsic images [75, 42, 72], classifier guidance [16] and cross-modal data [12, 69]. Notably, training-free guidance methods [43, 20, 3, 4] have enabled image editing capabilities without specific training, forming the basis of our approach.
@@
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for style transfer [25, 41] or harmonization [51, 46, 35, 61] do not prioritize physically accurate shading of the objects. Recently, [47] utilized a single-view multi-illumination dataset [44] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [65, 11] for relighting indoor scenes. IC-Light [73] enforces consistency in appearance to relight portraits and various objects. Some methods specialize in outdoor scenes, using geometric priors like depth [27] or normalized coordinates [32] or both [38] to improve global lighting estimation. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13]. ZeroComp [75] learns to composite virtual objects in a zero-shot fashion, but does not offer lighting control. RGBX [72] uses diffusion models for material- and lighting-aware composition with text-guided control. Neural Gaffer [30] and DiLightNet [71] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. We bridge this gap by introducing a shadow guidance for object relighting.
@@
Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 37, 76], though they generally lack controllability. Parametric models offer more control [23, 63, 15], but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
@@
Shadow generation Generative models for rendering realistic shadows [74, 39, 29, 40], or harmonizing rendered shadows with the image [62] have been proposed. Other methods enable controllable shadows for 3D [55] and 2D objects [56, 57]. Our approach can leverage existing shadow generation methods and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
@@
Image intrinsics Decomposing images into albedo and shading has long been studied [7, 34, 45, 64, 13]. Many works have also focused on estimating scene geometry, for example, depth [49, 8, 70, 31, 66, 67] and normals [18, 6, 50, 5, 68]. Recently, approaches such as [37, 76] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [10, 17, 31, 68, 2, 42]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
@@ 3 SpotLight
Diffusion-based neural renderers are currently limited by their inability to directly control shading, especially for specific objects as they are generally conditioned on albedo, normals, depth, and global shading. SpotLight addresses this limitation by enabling control over a single object lighting effects by defining a coarse shadow. Leveraging priors from pre-trained models and the diffusion framework, we introduce a method capable of harmonizing the object realistically from a provided shadow region. Fig. 2 presents an overview of the method, with each component described in detail below.
3.1 Rendering from intrinsics
Our method leverages a pre-trained diffusion-based neural renderer (e.g., ZeroComp [75], RGBX [72]) which accepts a multi-channel input of intrinsic maps —typically: albedo, shading, surface normals and depth, although other maps such as roughness and metallic can also be used—and outputs an RGB image consistent with these inputs. Our approach also requires shadow and object masks, and resp., where 1 indicates if the pixel is in shadow/object, and 0 otherwise. The discussion below assumes is based on latent diffusion [52], which performs the diffusion process in a low resolution latent space defined by a VAE encoder .
3.2 Blending shadows with the background
Provided a user defined shadow mask (), we first need to ensure that the diffusion model will properly blend the shadow in its final render. To achieve a perceptually pleasing shadow, we take inspiration from Blended Latent Diffusion [4] to progressively blend the noisy latents at time with a noised VAE-encoded image which contains the shadows composited over the image, . In practice, we obtain by compositing the object albedo and the shadow on the background. The update made at each timestep to incorporate the desired shadow latents is therefore
where is the shadow latent weight and is the shadow mask, first dilated by a square kernel to maintain the desired shadow softness on the edges, then bilinearly downsampled to the latent resolution (). Noise is added to the encoded composite using the DDIM [59] noise scheduler. Different from [4], we do not use the slow test-time decoder optimization nor the progressive mask shrinking steps.
3.3 Enhancing lighting control
Simply running a neural renderer conditioned on the desired shadow direction yields perceptible, but subtle visual changes in the generated object (as will be shown in sec. 4.6). We propose to amplify the effect by running two parallel branches of the model: one which receives the shadow in the desired direction (positive branch), while the other has the shadow in the opposite direction (negative branch), obtained by shifting the light azimuth angle by . Although the opposite light direction works best for the negative branch, we found that casting no shadow also provides a good negative sample. We obtain the outputs of the diffusion renderer, which are typically -predictions [54], that is, a linear combination of the noise and the estimated image. We then use classifier-free guidance to amplify the effect of the positive shadow on the lighting on the object:
where is the guidance scale, is the object mask bilinearly downsampled to the latent resolution (). After applying eq. 2, we run a regular DDIM sampling step to obtain the latents .
3.4 Diffusion-based neural renderer backbone
Since SpotLight only applies changes to the latent space of a diffusion model, it can readily be applied to pre-trained diffusion-based neural renderers without any finetuning.
ZeroComp [75] provides a maskable shading map as input to a ControlNet-based neural renderer. Here, we use their approach directly and mask out the shading on both the object and the shadow region.
RGBX [72] provides a checkpoint of their XRGB model finetuned for inpainting rectangular masks. We mask out the bounding box including the object and shadow region.
At each denoising step, the denoising network is ran to obtain
where represents the network -prediction, represents the composited intrinsics. For ZeroComp, those intrinsics correspond to albedo, normals, depth, and partial shading and for RGBX, we use the albedo, normals, metallic, roughness and masked image.
3.5 Obtaining the final render
After having performed the denosing steps, we obtain an image by decoding the latents into by using the VAE decoder. The final result is obtained with the background preservation strategy proposed in [75], which obtains a shadow matte by computing the ratio of the image obtained with and without shadow guidance. This ensures the diffusion process does not affect the background image.
@@ 4 Evaluation on 3D object compositing
In this section, we quantitatively evaluate SpotLight against other methods that support lighting control on composition of 3D objects in 2D images. Additionally, we conduct a user study to assess human perception, comparing its performance with these competing approaches.
@@ 4.1 Evaluation dataset
We start from the evaluation dataset proposed in [75], where high-quality 3D models from the Amazon Berkeley Objects dataset [14] are inserted into background images cropped with a 50° field of view from the Laval Indoor HDR dataset [22]. The authors of [75] graciously provided us with reproducible Blender scene information, which we extend in two ways. First, we improve the shadow catcher using back-projected background depth estimated using Depth Anything V2 [67]. Second, object intrinsics are rendered with anti-aliased alpha mattes, allowing for more realistic compositing results. We use a depth-warping strategy on the HDR panorama as in [75] to account for spatially varying lighting around the object position. The dataset is rendered using the physically-based Cycles renderer with 2048 samples per pixel, a noise threshold of 0.01, and denoised using OpenImageDenoise [1]. We produce two versions of the dataset, each intended for different purposes.
“Reference-based” dataset This version identifies the dominant light source in the scene, extracted from the ground-truth warped panorama using the light extraction algorithm from [23]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3. We filter out samples where the background depth estimation failed. This data set is intended for quantitative evaluation (sec. 4.4), allowing the simulated ground truth of the inserted object to be used to compute metrics.
“User-controlled” dataset This second dataset simulates a scenario where a user specifies a desired light position. In this setup, five lighting directions are defined from left to right in increments in azimuth, with the light source passing behind the object to ensure the shadow remains visible. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,050 images. This dataset will be used in the user study (sec. 4.5) to assess the perceptual accuracy of renders in edited lighting scenarios.
| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) | Simulated GT |
@@ 4.2 SpotLight specifics
As mentioned in sec. 3.4, our proposed SpotLight readily applies to existing diffusion-based neural renderers. Here, we experiment with two approaches from the recent literature: ZeroComp [75] and RGBX [72].
ZeroComp backbone We train ZeroComp following the procedure outlined in [75], and use the same conditioned intrinsic estimators at inference time: StableNormal [68] for normals, IID [33] for albedo, and ZoeDepth [8] for depth. The masked shading in the background is obtained by dividing the image by the albedo and applying the same masking strategy. Given its superior results, we adopt this method as our default model for 3D object compositing.
RGBX backbone We use the RGBX model to extract intrinsic properties (normals, metallic, roughness, and albedo). We observe that results are generally overly bright, likely due to a domain gap between the training and evaluation images. To address this, we re-expose the background by a factor of 2 before feeding it to the network, then divide the output by 2, in linear space. We apply the same background preservation strategy as in ZeroComp.
Light control SpotLight requires a coarse shadow for guidance. Therefore, we use a real-time rasterization renderer (Blender’s EEVEE) to compute a simple shadow of the 3D object. This shadow is generated from a simple parametric distant sphere light and projected onto the mesh constructed from the estimated depth map of the background.

### v2

@@ Abstract
Recent work has shown that diffusion models can serve as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. However, unlike typical physics-based renderers, these neural rendering engines are limited by the lack of manual control over the lighting, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise lighting control can be achieved for object relighting simply by providing a coarse shadow of the object. Indeed, we show that injecting only the desired shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting. We also demonstrate other applications, such as hand-scribbling shadows and full-image relighting, demonstrating its versatility.
Background image | Previous method | SpotLight enables shadow-guided lighting control |
@@ 1 Introduction
Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process. When one is given a photograph, however, no such control is possible. Relighting then becomes a hard task because of the complex interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture. The recent rise of powerful image generation models (e.g., [54]) has enabled exciting possibilities for image relighting [31, 76].
@@
This paper focuses on providing lighting control to recent diffusion image relighting frameworks, where one wishes to modify the appearance of a single object. This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image. In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
We present SpotLight, a method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [74, 29]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer. Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it to relight the object accordingly. We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object as it moves around it.
Fig. 1 shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [19] according to the desired lighting direction. From this, SpotLight realistically blends the input guiding shadow within the background, and relights the virtual object to match the desired lighting. To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
@@
Our core technical contribution lies in our shadow guidance approach, which refines the guiding shadow input while realistically relighting the virtual object. Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control. This builds upon existing object compositing networks [78, 75] that operate on intrinsic images along with diffusion models’ learned priors to render realistic outputs. We demonstrate that SpotLight’s guidance mechanism handles complex scenarios and maintains physical plausibility without requiring manual intervention nor any additional training.
@@
In summary, our contributions are as follows. First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring additional training. We demonstrate superior quality, both perceptually through user studies and quantitatively, compared to competitive specialized models. Finally, we release an evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
@@ 2 Related work
Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [59, 27, 60, 54]. These models offer control mechanisms for various modalities, such as text [50, 54], image content [55, 21], intrinsic images [78, 43, 75, 39], classifier guidance [16] and cross-modal data [12, 72]. Notably, training-free guidance methods [44, 20, 3, 4, 26] have enabled image editing capabilities without specific training, forming the basis of our approach.
@@
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for harmonization [53, 48, 34, 62] do not prioritize physically accurate shading of the objects. Recently, [49] utilized a single-view multi-illumination dataset [45] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [67, 11] for relighting indoor scenes. IC-Light [76] enforces consistency in appearance to relight portraits and various objects. LumiNet [68] and ScribbleLight [14] achieve image relighting through illumination transfer or user-defined scribbles, respectively. Some methods specialize in outdoor scenes, using geometric priors like depth [25] or normalized coordinates [31] or both [40]. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13]. Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [78, 75, 39], which can be leveraged to achieve zero-shot object composition. However, these methods have limited lighting control. ZeroComp [78] provides no lighting control, RGBX [75] can only provide lighting control through a text prompt, and DiffusionRenderer [39] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set. Other diffusion-based approaches, such as Neural Gaffer [29], DiLightNet [74] and IllumiNeRF [79] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. We bridge this gap by introducing a shadow-guided strategy for object relighting that leverages existing models without any additional training.
@@
Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 36, 80, 47, 38], though they generally lack controllability. Parametric models [23, 65, 15] or GAN inversion methods [64] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
@@
Shadow generation Generative models for rendering realistic shadows [77, 41, 28, 42], or harmonizing rendered shadows with the image [63] have been proposed. Other methods enable controllable shadows for 3D [56] and 2D objects [57, 58]. Our approach can leverage existing shadow generation methods (e.g., [58], see supp.) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
@@
Image intrinsics Decomposing images into albedo and shading has long been studied [7, 33, 46, 66, 13]. Numerous works have also focused on estimating scene geometry, for example, depth [51, 8, 73, 30, 69, 70] and normals [18, 6, 52, 5, 71]. Recently, approaches such as [36, 80] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [10, 17, 30, 71, 2, 43]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
3 Background: diffusion renderers
Recent work [75, 78] has shown that diffusion models can be adapted to generate images from intrinsic maps including materials (e.g., albedo, roughness, metallic), geometry (e.g., surface normals and depth), and shading. This capability was built using a ControlNet with ZeroComp [78] or full finetuning with RGBX [75]. We refer to this new class of conditional generation models as “diffusion renderers”.
In these models, the denoising network is ran at each denoising step to obtain
where represents the network -prediction, the noisy latents, and the conditioning intrinsic maps.
It has been shown that these methods can realistically composite virtual objects into images in a zero-shot fashion, despite not being trained for this task. Provided a mask of the virtual object and its intrinsic maps (rendered with graphics shaders), the object can be trivially composited in most intrinsic maps of the background image (estimated with trained estimators). Since computing the virtual object shading requires complex lighting simulations, diffusion approaches estimate the shading by providing a partial shading map [78], or train inpainting models [75]. However, unlike traditional rendering, these methods offer no control over the lighting conditions.
In this work, we use a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, analogous to a spotlight moving around the object. We now describe our approach in detail.
@@ 4 SpotLight
We employ a user-provided, guiding shadow to steer the pre-trained diffusion process, thereby enforcing matching illumination on the inserted object. Our key insight is that even an approximate shadow encapsulates important lighting information. By integrating this cue early in the diffusion process, we simultaneously refine both the composite appearance and its shadow, yielding a natural blend with the scene.
4.1 Approach overview
Our approach requires two inputs: an object mask and a user-specified guiding shadow . This shadow, computed via fast rasterization or specified through hand-drawn scribbles, is used to guide the latent diffusion process. Without retraining the pre-trained renderer , we steer its generative process by incorporating the guiding shadow into the latent space.
4.2 Blending shadows
To fuse the guiding shadow with the latent representation, we update the noisy latent code at each diffusion timestep, similar to [4]. In our framework, the updated latent
serves as a soft constraint that steers the process toward a shadow-consistent output. Here, is the composite of the object albedo, the target shadow, and the background, and controls the strength of the guiding shadow. The latent shadow mask is obtained by downsampling the target shadow to the latent resolution (see supp. for details).
4.3 Shading the object via dual-branch guidance
A key element of our approach is to amplify the impact of the guiding shadow on the object shading through classifier-free guidance (CFG) [26]. We process two parallel diffusion branches: a positive branch that incorporates the desired shadow direction, and a negative branch that suppresses an undesired shadow direction (e.g., via an opposing light direction). Fig. 2 shows our dual-branch guidance mechanism. Classifier-free guidance then combines these branches:
where sets the guidance scale and is the object mask downsampled to the latent resolution. This dual-branch strategy reinforces the intended lighting on the object while naturally blending it with the background.
4.4 Final image synthesis
After the guided diffusion process, the resulting latent is decoded with the VAE decoder. The background preservation strategy of [78] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
@@ 5 Evaluation on 3D object compositing
Evaluating composition methods where lighting conditions can be controlled is challenging, as the quality of the results must be assessed across multiple lighting condition for the same scene. To address this, we build over the evaluation setup of [78] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas. We evaluate two scenarios: (1) “reference-based”, which preserves the original scene lighting to assess the ability of methods to produce images close to the ground truth, and (2) “user-controlled”, where the original lighting is modified (e.g., dominant light rotation) to simulate user-specified lighting conditions, requiring methods to adapt while maintaining a perceptually pleasing render. In this section, we describe how these scenarios are used for quantitative evaluation and a user study to assess the realism and controllability of the methods.
@@ 5.1 Evaluation dataset
We create datasets for the two scenarios using the Blender scene information graciously provided by the authors of [78]. We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [70]. We also improve the ground truth intrinsic maps by providing correct antialiased versions.
“Reference-based” scenario To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [78] to ensure lighting fidelity around the inserted object. We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [19]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3, in the “ground truth” column. This dataset is intended for quantitative evaluation (sec. 5.4).
“User-controlled” scenario This dataset simulates a scenario where a user specifies a desired light position. 8 lighting directions are defined, spaced at increments in azimuth around the object, and used to create the guiding shadows, again using blender’s EEVEE engine. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,680 images. This dataset is utilized in the user study (sec. 5.5) to evaluate the perceptual accuracy of renders under modified lighting conditions.
@@ 5.2 SpotLight specifics
As mentioned in sec. 3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [78] or RGBX [75]. Here, we provide details on ZeroComp, see supp. for RGBX.
We use the checkpoint trained on the Openrooms dataset [37] provided by the authors. The masked shading map is obtained by dividing the background image by the albedo, and by masking out the shading on both the object and shadow regions. Given its superior results, we adopt this method as our default model for 3D object compositing. To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [78], namely StableNormal [71] for normals, IID [32] for albedo, and ZoeDepth [8] for depth.
Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
Light-conditioned methods | |||||
DiLightNet [74] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
Neural Gaffer [29] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
Shadow-conditioned methods | |||||
IC-Light [76] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
ZeroComp+SDEdit [44] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
SpotLight | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |